Properties of robust vertex fitting algorithms at high luminosity HEP experiments

J. D’Hondt^{a}, R. Frühwirth^{b}, S. Tavernier^{a}, P. Vanlaer^{c} and W. Waltenberger^{b}

(a) Vrije Universiteit Brussel,Pleinlaan 2, B-1050 Brussels, Belgium

(b) Universität Wien, Nikolsdorfer Gasse 18, A-1050 Vienna, Austria

(c) Université Libre de Bruxelles, Boulevard du Triomphe, B-1050 Brussels, Belgium

Abstract

Robust vertex fitting algorithms are expected to improve the knowledge of the vertex position and of its uncertainty, in the presence of mis-measured or misassociated tracks. Such contaminations are likely to happen in high luminosity HEP experiments. We have performed a simulation study of the sensitivity of two types of robust algorithms : a trimmed least-squares estimator and an adaptive estimator. Their statistical properties are studied as a function of the source and the level of contamination, and compared to the results obtained with classical least-squares estimators.

(Contribution for the BPS International Scientific Meeting 2004)